# ECLARE: Efficient cross-planar learning for anisotropic resolution enhancement

### v1

@@ 3.2 Approximating the imaging process
The imaging process described with Eq. 1 comprises blurring with $h$ and subsampling by $r$. In this subsection we discuss how we approximate both of these. In general, the slice selection profile $h$ is proprietary to the scanner manufacturer and cannot be derived precisely. We leverage ESPRESO (Han et al., 2023) to estimate $h$, shown in Fig. 1 (A). Briefly, ESPRESO optimizes a neural network to produce a fixed-size discrete $h$ by adversarial training. It trains on a single LR anisotropic volume by assuming that degraded in-plane patches and through-plane patches are indistinguishable from one another for an appropriate patch size. That is, ESPRESO assumes in-plane and through-plane patches come from the same distribution. Thus, in-plane image patches are extracted, degraded by $h$, and judged through a discriminator against through-plane patches. An optimal $h$ yields degraded in-plane image patches that are indistinguishable from true through-plane image patches. Every component is now present for Eq. 1: $h$ is approximated by ESPRESO and $r$ is known since we have a HR slice separation and LR slice separation.
Figure 2: A 1D demonstration of the differences between interpolation implementations. The input has 6 pixels and its FOV center is located at 2.5, indicated by the red dashed line. Its sample spacing is 1, indicated by the purple arrow. Its extents occur at -0.5 and 5.5, indicated by the blue dashed line. In this demonstration, the desired sample spacing for upsampling is 0.7, and the desired sample spacing for downsampling is 1.429. The implementation scipy.zoom (Virtanen et al., 2020) maintains the FOV center, but changes the sample spacing to 0.625 and 1.667 for upsampling and downsampling, respectively. The implementation torch.nn.functional.interpolate (Paszke et al., 2019) anchors the FOV extent to start at 0.5, shifting the center from the desired FOV center. Our proposed method (Sec. 3.3) maintains both the sample spacing and the FOV center.
@@ 3.3 FOV-aware interpolation
While interpolation is a routine operation in digital signal processing, the exact implementation in software has several freedoms. ECLARE requires both downsampling (for training data simulation) and upsampling (for inference) and we would like both to be stable with respect to the FOV. We consider three conditions that are desirable to meet with respect to the FOV of the signal:
@@
    the extents of the FOV should be maintained.
Interpolation algorithms cannot meet all of these conditions in general since only a discrete number of samples is permitted. However, since it is possible to satisfy two of these conditions, a compromise on the third is necessary. We quickly review popular modern implementations. The implementation scipy.ndimage.zoom (Virtanen et al., 2020) meets conditions 1) and 3). It adjusts the sample spacing to preserve the FOV center and extents. The implementation torch.nn.functional.interpolate (Paszke et al., 2019) meets only condition 2). It preserves the sample spacing, but neither the FOV center nor the extents. Instead, it anchors the beginning of the FOV to match the input. In our work, we place more importance on the center of the FOV and the sample spacing and therefore desire conditions 1) and 2). We summarize the interpolation method proposed in (Han, 2022) that maintains the FOV center and sample spacing while sacrificing the extents of the FOV. A comparison of the three implementations is shown in Fig. 2.
As in Han (2022), we introduce a fixed coordinate system for the signal before and after interpolation. Let the position of the first pixel be centered at zero, let $N$ be the number of pixels before interpolation, and let the spacing before interpolation be one. The width of each pixel equals the spacing; before interpolation it is therefore also one. The first pixel $p_{0}$ is centered at zero and the last pixel $p_{N-1}$ is centered at $N-1$. Since each pixel has a width of one, $p_{0}$ extends from $-0.5$ to $0.5$ and $p_{N-1}$ extends from $N-1.5$ to $N-0.5$. Therefore the entire image extends from $-0.5$ to $N-0.5$.
@@
This allows us to notate the first pixel after interpolation as $p^{\prime}_{0}$ and the last pixel as $p^{\prime}_{N^{\prime}-1}$.
@@
Also, since we know the new spacing $d$ and the new number of pixels $N^{\prime}$, we can calculate the position of any pixel $p_{i}$, $i=0,1,2,\ldots,N^{\prime}-1$ in terms of the new starting pixel:
@@
Finally, we can calculate the extents of the entire image after interpolation. The first pixel extends from $p^{\prime}_{0}-\frac{d}{2}$ to $p^{\prime}_{0}+\frac{d}{2}$ and the last pixel extends from $p^{\prime}_{N-1}-\frac{d}{2}$ to $p^{\prime}_{N-1}+\frac{d}{2}$. Thus, the image extents are $p^{\prime}_{0}-\frac{d}{2}$ to $p^{\prime}_{N-1}+\frac{d}{2}$.
The result of this method yields an FOV-aware interpolation implementation that maintains the FOV center and guarantees the desired pixel spacing. This is true in both upsampling and downsampling scenarios and is demonstrated in Fig. 2.
@@ 3.4 Training a deep learning model with paired data
We now develop a relationship between in-plane and through-plane data and train a SR network. The supervised learning framework in SR requires paired LR and HR data. We extract HR 2D patches from in-plane and produce blurred 2D patches using Eq. 1 with the learned $h$ from ESPRESO. These blurred 2D patches are then downsampled using the FOV-aware interpolation described in Sec. 3.3. Thus, each LR patch has the appropriate simulated slice thickness and separation while additionally being FOV-centered according to the HR patch.
Figure 3: A diagram of our proposed architecture, based on WDSR (Yu et al., 2018). The low resolution image $y$ is first processed by a convolution to produce a feature map with 256 channels. Simultaneously, it is upsampled to produce an interpolated image to be used later as a skip connection. It then passes through several WDSR blocks. A final convolution is done to produce the correct number of channels for pixel shuffle. Finally, the result is downsampled and added to the skip connection to yield the output. In our implementation, we set $C$ to 256 and use 16 WDSR blocks.
A deep convolutional neural network model learns the mapping between the paired patches, as shown in Fig. 1 (B). We use an architecture based on WDSR (Yu et al., 2018) after empirically evaluating other architectures such as EDSR (Lim et al., 2017) and RCAN (Zhang et al., 2018). The WDSR architecture, designed for efficiency, achieved the same quantitative metrics as other architectures but trained in less time. However, the WDSR architecture itself uses pixel shuffle (Shi et al., 2016), which requires integer upsampling factors, to perform the final upsampling to the desired image size. In general, though, $r$ may be non-integer. To address this, we rewrite $r$ as
@@ 3.6 Implementation details
ECLARE is implemented in PyTorch and we train the model from scratch with each new multi-slice 2D MR image. We optimize the parameters of the network by minimizing mean squared error using a total of $1,000,000$ patches with a batch size of $128$ patches. The training patches are of size $\lceil 8r\rfloor\times 8$. We use a $21$-tap filter to represent $h$, which follows the standard implementation of ESPRESO. We used the Adam (Kingma and Ba, 2014) optimizer, a learning rate of $10^{-3}$, the one-cycle learning rate scheduler (Smith and Topin, 2019), and PyTorch’s automatic mixed precision feature. Our implementation uses $16$ residual blocks of $256$ channels. With these hyperparameter choices, on an NVIDIA V100 the total running time, including self-training and inference, is just under 5 minutes. The memory usage is 1.16 GB on the GPU. Our code is open-source and is available at https://www.github.com/sremedios/eclare.
@@ 4 Experiments and Results
To determine the efficacy of ECLARE, we designed two simulation experiments and one experiment for real-world evaluation.
@@ 4.1 Data and evaluation metrics
The simulation experiments began with HR isotropic 3D-acquired image volumes of resolution 1×1×1 $1\times 1\times 1\text{~{}mm}^{3}$. These images were then degraded with Eq. 1, setting $h$ to the shape of the slice excitation profile created by the Shinnar-Le Roux (SLR) pulse design algorithm (Pauly et al., 1991; Shinnar et al., 1989), $s$ to the desired slice spacing (described below), and $x$ to a 3D-acquired isotropic volume. The implementation of the SLR algorithm we use is from the sigpy library (Martin et al., 2020).
@@
We compare ECLARE to bicubic interpolation, SynthSR (Iglesias et al., 2023) (Sec. 2.3), single-contrast INR (McGinnis et al., 2023) (Sec. 2.2), and SMORE (Zhao et al., 2020). Since its output is always a T1-w MPRAGE, we do not evaluate SynthSR on the FLAIR SR task.
To highlight the advantages of the independence from external training data, we also show results on a variety of real-world data, as follows,
  1. 1.
    a $5\|1$ T2-w FLAIR scan of a human head from the ultra low-field (ULF) ($0.3$T) M4Raw dataset (Lyu et al., 2023) with an SR factor of $r=6$;
  2. 2.
[…]
  3. 3.
[…]
  4. 4.
[…]
For these real-world data, as no ground truth images are available, results are purely qualitative. Additionally, since none of the images are T1-w MPRAGE human heads, SynthSR was excluded from this evaluation; thus the comparison methods included bicubic interpolation, single-contrast INR, SMORE, and ECLARE.
Since we have ground truth (GT) images for simulated data, we can use reference-based image quality metrics to evaluate SR performance. We use peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) with the implementations provided by the scikit-image Python library (Van der Walt et al., 2014). We additionally evaluate the utility of the super-resolved images by segmentation downstream tasks. We use the consistency Dice similarity coefficient (cDSC), which we define as the Dice similarity coefficient (DSC) between an automated algorithm’s performance on the HR image volume and the super-resolved image volume. This is to differentiate cDSC from DSC which commonly is used between an algorithm-generated label and a human-generated label.
@@ 4.2 Healthy T1-w simulations
We calculated cDSC on the SLANT (Huo et al., 2019) and SynthSeg (Billot et al., 2023) estimates of the OASIS-3 dataset. Since SLANT produces many small labels, we simplify them down to $9$ labels as in our previous work (Remedios et al., 2023): ventricles, cerebellar gray matter (GM), cerebellar white matter (WM), cerebral GM, cerebral WM, caudate, putamen, thalamus, and brainstem. SynthSeg is a generalized automated segmentation technique designed to be robust to a variety of image appearances. It directly produces volume estimates of similar regions of interest, so we include the following comparable regions: ventricles, cerebral white matter, cerebral gray matter, cerebellar white matter, cerebellar grey matter, thalamus, putamen, caudate, and brainstem.
Figure 4: A $5\|1.5$ sagittal slice from a representative subject from the OASIS-3 dataset. The first row shows the image domain with PSNR and SSIM for the entire image volume overlaid. The second row shows the simplified SLANT labels, with an overlay showing the mean cDSC across all labels for the entire image volume. The third row shows the log Fourier magnitude of the corresponding slice. The fourth row shows the absolute difference between the log Fourier magnitudes of the SR method and the ground truth.
@@
On the OASIS-3 dataset, we show qualitative results from a representative subject at $5\|1.5$ resolution in Fig. 4. The image domain is shown alongside the SLANT result labels, the log Fourier magnitude, and the error in log Fourier magnitude across the methods. In the image domain, it is apparent that our proposed ECLARE method is visually closer to the ground truth (GT) in both appearance and contrast. Viewing Fig. 4 column-by-column, in the image domain bicubic interpolation does not address the extreme aliasing, but is relatively accurate with respect to the SLANT labels. In the Fourier domain, bicubic interpolation does not recover high-frequency information, and thus the error shown in the Fourier magnitude is only in high-frequency areas. SynthSR generates realistic brain and head appearance, but mismatches the contrast and hallucinates some structures such as the cortical sulci that can be seen in the SLANT labels row. Since SynthSR regenerates the entire image, we see a loss of resolution in the Fourier domain along the anterior-posterior axis as well as yielding an error at zero frequency (i.e., a change in DC). Single-contrast INR does not generate accurate results in the image domain, and actually synthesizes a different brain stem appearance from the other methods, as shown in the green brainstem label in the second row. However, single-contrast INR does generate the most high-frequency components in Fourier space, though it also makes errors in low-frequency regions. The image domain results for SMORE highlight an “over-sharpening” effect with hyperintense regions near the skull and superior to the ventricles and the bottom part of the brainstem. As shown in the log Fourier magnitude error row of Fig. 4, both SMORE and ECLARE makes fewer errors in the low-frequency regions of Fourier space. However, our proposed ECLARE improves upon the results of SMORE, yielding the most visually accurate image of the group and the least low-frequency errors in Fourier space, even compared to bicubic interpolation. The SLANT segmentation results are still not ideal for any method, and some areas of cortical gray matter are missed by all methods.
@@
We show the PSNR and SSIM for the five methods in Fig. 5. Across all resolutions, our proposed ECLARE outperforms all other methods in both metrics except for the resolution $3\|0$, where SMORE is best. SMORE’s superiority at $3\|0$ is likely due to two reasons: 1) SMORE does not have a model of slice gap, and therefore will have peak performance when no gap is present in the data; and 2) SMORE relies on pre-trained weights and was primarily tested and evaluated at data of resolutions $2\|0$ and $3\|0$.
@@
In Fig. 6, we show the cDSC for the nine regions of interest calculated by SLANT (Huo et al., 2019). Overall we see similar trends as in Fig. 5, with ECLARE performing best at all resolutions except at $3\|0$, where SMORE performs best. An exception to this is the cerebellar white matter, where at nearly every resolution SMORE is most performant. An interesting observation is how robust SLANT is to interpolation artifacts; bicubic interpolation outperforms the other two super-resolution methods consistently across all resolutions and structures.
@@
We also used SynthSeg (Billot et al., 2023) as a downstream task. For all subjects, we fit a linear regression model to the sets of estimated volumes against the GT volumes per region per resolution where the dependent variable was the GT volume and the independent variable was the SR volume. The fit of the line is quantified by the $R^{2}$ coefficient of determination, which we plot in Fig. 7. Some regions (ventricles, cerebral white matter, brainstem) are consistently well-segmented at all resolutions across all SR methods, but others show clear winners. ECLARE outperforms all methods consistently, visualized by the enclosure of other methods on the radar plots.

### v2

@@ 3.2 Approximating the imaging process
The imaging process described by Eq. 1 comprises blurring with $h$ and subsampling by $r$. The subsampling factor $r$ is known since the HR and LR slice separations are each known, so the remaining task is to determine the slice selection profile $h$. In general, $h$ is a product of a real-world implementation that is subject to imperfections in the instrumentation, and hence cannot be derived precisely. However, since our intent is to use HR in-plane patches as the target distribution, estimating $h$ directly is not necessary. Instead, we leverage ESPRESO [30] to estimate a relative slice selection profile $\tilde{h}$, shown in Fig. 1 (A). The desired relative slice selection profile $\tilde{h}$ blurs data from the in-plane slices such that the distribution matches the LR through-plane data. To estimate $\tilde{h}$, ESPRESO optimizes a neural network to produce a fixed-size discrete profile through adversarial training. It trains on a single LR anisotropic volume by assuming that degraded in-plane and through-plane patches are indistinguishable in distribution for an appropriate patch size. Briefly, training consists of extracting in-plane image patches, degrading by $\tilde{h}$, comparing to through-plane patches, and backpropagating the loss through a discriminator. The optimal $\tilde{h}$ yields degraded in-plane image patches that are indistinguishable in distribution from true through-plane image patches. We use this estimated $\tilde{h}$ in lieu of $h$ for Eq. 1 hereafter.
@@ 3.3 FOV-aware interpolation
While interpolation is routine in digital signal processing, the exact implementation has several freedoms. ECLARE requires both downsampling (for training data simulation) and upsampling (for inference) and we would like both to be stable with respect to the FOV. We consider three conditions that are desirable to meet with respect to the FOV of the signal:
@@
    the locations of boundary samples should be maintained.
Interpolation algorithms cannot meet all of these conditions in general since only a discrete number of samples is permitted. However, since it is possible to satisfy two of these conditions, a compromise on the third is necessary. We quickly review popular modern implementations. The method scipy.ndimage.zoom [38] meets conditions 1) and 3). It adjusts the sample spacing to preserve the FOV center while locking the locations of boundary samples. The method torch.nn.functional.interpolate [39] meets only condition 2). It preserves the sample spacing, but neither the FOV center nor the boundary sample locations. Instead, it anchors the extent on the first sample. We place more importance on the center of the FOV and the sample spacing and therefore desire conditions 1) and 2). We summarize the interpolation method proposed in [52] that maintains the FOV center and sample spacing while sacrificing the precise location of the boundary samples. A comparison of the three implementations is shown in Fig. 2.
Figure 4: A $5\|1.5$ sagittal slice from a representative subject from the OASIS-3 dataset. The first row shows the image domain with PSNR and SSIM for the entire image volume overlaid. The second row shows the simplified SLANT labels, with an overlay showing the mean cDSC across all labels for the entire image volume.
As in [52], we introduce a fixed coordinate system for the signal before and after interpolation. Let the position of the first pixel be centered at zero, let $N$ be the number of pixels before interpolation, and let the spacing before interpolation be one. The width of each pixel is equal to the spacing; before interpolation it is therefore also one. The first pixel $p_{0}$ is centered at zero and the last pixel $p_{N-1}$ is centered at $N-1$. Since each pixel has a width of one, $p_{0}$ extends from $-0.5$ to $0.5$ and $p_{N-1}$ extends from $N-1.5$ to $N-0.5$. Therefore the entire image extends from $-0.5$ to $N-0.5$.
@@
Accordingly, we notate the first pixel after interpolation as $p^{\prime}_{0}$ and the last pixel as $p^{\prime}_{N^{\prime}-1}$.
@@
Also, since we know the new spacing $d$ and the new number of pixels $N^{\prime}$, we can calculate the position of any pixel $p_{i}^{\prime}$, $i=0,1,2,\ldots,N^{\prime}-1$ in terms of the new starting pixel:
@@
Finally, we can calculate the new extents of the entire image after interpolation. The first pixel extends from $p^{\prime}_{0}-\frac{d}{2}$ to $p^{\prime}_{0}+\frac{d}{2}$ and the last pixel extends from $p^{\prime}_{N-1}-\frac{d}{2}$ to $p^{\prime}_{N-1}+\frac{d}{2}$. Thus, the image extents are $p^{\prime}_{0}-\frac{d}{2}$ to $p^{\prime}_{N-1}+\frac{d}{2}$.
The result of this method yields an FOV-aware interpolation implementation that maintains the FOV center and guarantees the desired pixel spacing. This is true in both upsampling and downsampling scenarios as demonstrated in Fig. 2.
@@ 3.4 Training a deep learning model with paired data
We now develop a relationship between in-plane and through-plane data and train a SR network. Paired LR and HR data are necessary for supervised learning. We extract HR 2D patches from in-plane and produce blurred 2D patches using Eq. 1 with the learned $\tilde{h}$ from ESPRESO. These blurred 2D patches are then downsampled using the FOV-aware interpolation described in Sec. 3.3. Thus, each LR patch has the appropriate simulated slice thickness, separation, and FOV.
A deep convolutional neural network model $f_{\theta}$ learns the mapping between the paired patches, as shown in Fig. 1 (B). We use an architecture based on WDSR [37] after empirically evaluating other architectures such as EDSR [42] and RCAN [31]. The WDSR architecture, designed for efficiency, achieves the same quantitative metrics as other architectures but trains in less time. However, the WDSR architecture itself uses pixel shuffle [32], which requires integer upsampling factors, to perform the final upsampling to the desired image size. In general, though, the required upsampling ratio $r$ may be non-integer. To address this, we rewrite $r$ as
@@ 3.6 Implementation details
ECLARE is implemented in PyTorch and we train the model from scratch with each new multi-slice 2D MR image. We optimize the parameters of the network by minimizing mean squared error using a total of $1,000,000$ patches with a batch size of $128$ patches. The training patches are of size $\lceil 8r\rfloor\times 8$. We use a $21$-tap filter to represent $h$, which follows the standard implementation of ESPRESO. We used the Adam [53] optimizer, a learning rate of $10^{-3}$, the one-cycle learning rate scheduler [54], and PyTorch’s automatic mixed precision feature. Our implementation uses $16$ residual blocks of $256$ channels. With these hyperparameter choices, on an NVIDIA V100 the total running time, including self-training and inference, is just under 5 minutes. The memory usage is 1.16 GB on the GPU. Our code is open-source and is available at https://www.github.com/sremedios/eclare.
@@ 4.1 Data and evaluation metrics
Our experiments began with HR isotropic 3D-acquired image volumes of resolution 1×1×1 $1\times 1\times 1\text{~{}mm}^{3}$. These images were then degraded with Eq. 1, setting $h$ to the shape of the slice excitation profile created by the Shinnar-Le Roux (SLR) pulse design algorithm [50, 51], $s$ to the desired slice spacing (described below), and $x$ to a 3D-acquired isotropic volume. The implementation of the SLR algorithm we use is from the sigpy library [55].
@@
We compare ECLARE to bicubic interpolation, SynthSR [49], single-contrast INR [48], and SMORE [34]. Since its output is always a T1-w MPRAGE, we do not evaluate SynthSR on the FLAIR SR task. Other mainstream SR methods are not appropriate for comparison here due to the heterogeneities in data acquisition strategies for multi-slice 2D MRI. Many LR volumes have non-integer ratios to achieve isotropic resolution or have slice gap. Furthermore, there is a need to evaluate on arbitrary image contrasts without retraining for each subtle change in contrast. ECLARE and the methods we compare to are agnostic to these differences.
Since we have ground truth (GT) images for simulated data, we can use reference-based image quality metrics to evaluate SR performance. We use peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) with the implementations provided by the scikit-image Python library [58]. We additionally evaluate the utility of the super-resolved images by segmentation downstream tasks. We use the consistency Dice similarity coefficient (cDSC), which we define as the Dice similarity coefficient (DSC) between an automated algorithm’s performance on the HR image volume and the super-resolved image volume. This is to differentiate cDSC from DSC which commonly is used between an algorithm-generated label and a human-generated label.
@@ 4.2 Healthy T1-w simulations
We calculated cDSC on the SLANT [59] and SynthSeg [60] estimates of the OASIS-3 dataset. Since SLANT produces many small labels, we simplify them down to $9$ labels as in our previous work [36]: ventricles, cerebellar gray matter (GM), cerebellar white matter (WM), cerebral GM, cerebral WM, caudate, putamen, thalamus, and brainstem. SynthSeg is a generalized automated segmentation technique designed to be robust to a variety of image appearances. It directly produces volume estimates of similar regions of interest, so we include the following comparable regions: ventricles, cerebral white matter, cerebral gray matter, cerebellar white matter, cerebellar grey matter, thalamus, putamen, caudate, and brainstem.
@@
On the OASIS-3 dataset, we show qualitative results from a representative subject at $5\|1.5$ resolution in Fig. 4. The image domain is shown alongside the SLANT result labels across the methods. In the image domain, it is apparent that our proposed ECLARE method is visually closer to the ground truth (GT) in both appearance and contrast. Viewing Fig. 4 column-by-column, in the image domain bicubic interpolation does not address the extreme aliasing, but is relatively accurate with respect to the SLANT labels. SynthSR generates a realistic brain and head appearance, but mismatches the contrast and hallucinates some structures such as the cortical sulci that can be seen in the SLANT labels row. Single-contrast INR does not generate accurate results, synthesizing a different brain stem appearance from the other methods, as shown in the green brainstem label in the second row. The SMORE results highlight an “over-sharpening” effect with hyperintense regions near the skull and superior to the ventricles and the bottom part of the brainstem. ECLARE yields the most visually accurate image of the group in both the image as well as qualitative inspection of the segmentation. The SLANT segmentation results are still not ideal for any method, and some areas of cortical gray matter are missed by all methods.
@@
We show the PSNR and SSIM for the five methods in Fig. 5. Across all resolutions, our proposed ECLARE outperforms all other methods in both metrics except for the resolution $3\|0$, where SMORE is best.
@@
In Fig. 6, we show the cDSC for the nine regions of interest calculated by SLANT [59]. Overall we see similar trends as in Fig. 5, with ECLARE performing best at all resolutions except at $3\|0$, where SMORE performs best. An exception to this is the cerebellar white matter, where at nearly every resolution SMORE is most performant. An interesting observation is how robust SLANT is to interpolation artifacts; bicubic interpolation outperforms single-contrast INR and SynthSR consistently across all resolutions and structures.
@@
We also used SynthSeg [60] as a downstream task. For all subjects, we fit a linear regression model to the sets of estimated volumes against the GT volumes per region per resolution where the dependent variable was the GT volume and the independent variable was the SR volume. The fit of the line is quantified by the $R^{2}$ coefficient of determination, which we plot in Fig. 7. Some regions (ventricles, cerebral white matter, brainstem) are consistently well-segmented at all resolutions across all SR methods, but others show clear winners. ECLARE outperforms all methods consistently, visualized by the enclosure of other methods on the radar plots.